OpenStudy (perl):

A shipment of 12 microwave ovens contains 3 defective units. A vending company has ordered 4 of these units, and because all are packaged identically, the selection will be at random. What is the probability that at least 2 units are good?

OpenStudy (perl):

why isnt this question a binomial problem. oh because it is not with replacement

OpenStudy (callisto):

I'm not sure for this .... probability that at least 2 units are good = 1 - P(3 are bad) = 1- 3/12 x 2/12 x 1/12 x 9/12 = 1 - 1/384 I'm really not good at Probability

OpenStudy (perl):

ok i give answer

OpenStudy (perl):

P ( X >= 2) = P (x=2) + P(x=3) + p(x=4)

OpenStudy (perl):

P(x=2) = 9/12 *8/11*3/10*2/9 * 4!/(2!2!)

OpenStudy (perl):

hero, do you agree?

OpenStudy (anonymous):

Maybe this will help you: http://answers.yahoo.com/question/index?qid=20110224214712AAmLuiv

OpenStudy (anonymous):

easier to compute the probability that three are bad and subtract from one

OpenStudy (anonymous):

probability that in the shipment of 4 you get the 3 bad ones is \[\frac{\dbinom{9}{4}}{\dbinom{12}{4}}\]

OpenStudy (inkyvoyd):

1-(3/12)*(2/11)*(1/10)*(9/9)*4.

OpenStudy (inkyvoyd):

That's what I got.

OpenStudy (perl):

it says at least 2 , so P( X > = 2 ) = P ( X = 2) + P ( X = 3) + P ( X = 4)

OpenStudy (inkyvoyd):

perl, does what I say make sense?

OpenStudy (perl):

let me check

OpenStudy (perl):

@satellite that is wrong

OpenStudy (anonymous):

P ( x ≥ 2) = [ 9C2 * 3C2 + 9C3* 3C1+ 9C4 ] / 12C4 = 54/55

Directrix (directrix):

Probability of a defective oven = 3/12 = 1/4 = .25 Probability of a non-defective oven = .75 P( 2 Defects) = C(4,2) * (.25)^2 * (.75)^2 = .2109 P( 3 Defects) = C(4,3) * (.25)^3 * (.75)^1 = .0469 P(4 Defects) = C(4,4) * (.25)^4 * (.75)^0 = .0039 Probability of at least 2 defective ovens = .2109 + .0469 + .0039 = .2617 http://openstudy.com/study#/updates/4f7935efe4b0ddcbb89ea1e9

OpenStudy (perl):

binomial is with replacment, this problem is without replacement

OpenStudy (inkyvoyd):

@inkyvoyd is probably wrong.

Directrix (directrix):

A binomial model is characterized by trials which either end in success (heads) or failure (tails). These are sometimes called Bernoulli trials . Suppose we have n Bernoulli trials and p is the probability of success on a trial. Then this is a binomial model if 1. The Bernoulli trials are independent of one another. 2. The probability of success, p, remains the same from trial to trial. http://www.stat.wmich.edu/s160/book/node33.html

OpenStudy (perl):

this is no binomial problem